OpenStudy (anonymous):

How do you know which term goes first when solving for partial fractions, so you dont solve the a and b wrong? So if you have (x-3)(x+4) as the denominator, how do you know which term to make a and which to make b? Im assuming a wrong choice would reverse the values?

OpenStudy (amistre64):

it makes no difference

OpenStudy (anonymous):

easither one could go first but in the end youre going to use one of them first

OpenStudy (anonymous):

Ahhh cheers, dont worry, i got asked this question by a mate over facebook and it stumped me. Wasnt sure how to answer it

OpenStudy (amistre64):

spose you find that A = 1 and B = -2 as solutions in one setup such that you have: \[\frac{1}{x-3}+\frac{-2}{x+4}\]this is equal to\[\frac{-2}{x+4}+\frac{1}{x-3}\]by commutative property of addition

OpenStudy (anonymous):

But if you resolve the partial fractions it gives the same result anyway no matter what a/b term you have

OpenStudy (anonymous):

hello again

OpenStudy (anonymous):

yep lach

OpenStudy (anonymous):

Thanks very much guys. I feel very slapstick now :S

OpenStudy (anonymous):

this is what i was trying to say before, but i think i confused you as @amistre64 said it makes no difference but that is precisely why you have to write it down, to know which one you picked for A and B so you don't get confused at the end

OpenStudy (anonymous):

None the less thankyou all very much for the assistance! Hope you all have a good one!

OpenStudy (anonymous):

Ahhh I getcha now satellite, thanks very much mate!

OpenStudy (anonymous):

lol @Lachlan1996

OpenStudy (anonymous):

Hahaha, i feel bloody dopey now, but at least i figured it out ;) Thanks for the help everyone!!

OpenStudy (anonymous):

@Lachlan1996 youre from UK nice

OpenStudy (anonymous):

@SYNDICATE0612 hahahaha nahh mate, southern hemisphere ;) Proud bogan aussie haha

OpenStudy (anonymous):

nice @lachlan1996